OpenStudy (anonymous):

(3n+5)(2n-6)=0 Is this foil? How do I solve for n?

OpenStudy (anonymous):

yes use foil and u will get a quadratic equation. on solving the quadratic equation, you will get the ans.... if u need help first use foil and then let me know....

OpenStudy (anonymous):

Either 3n + 5 or 2n - 6 = 0 3n + 5 = 0 2n - 6 = 0 Solve each for n.